# On uniqueness in structured model learning

### v2

@@ Abstract
This paper addresses the problem of uniqueness in learning physical laws for systems of partial differential equations (PDEs). Contrary to most existing approaches, it considers a framework of structured model learning, where existing, approximately correct physical models are augmented with components that are learned from data. The main result of the paper is a uniqueness result that covers a large class of PDEs and a suitable class of neural networks used for approximating the unknown model components. The uniqueness result shows that, in the idealized setting of full, noiseless measurements, a unique identification of the unknown model components is possible as regularization-minimizing solution of the PDE system. Furthermore, the paper provides a convergence result showing that model components learned on the basis of incomplete, noisy measurements approximate the regularization-minimizing solution of the PDE system in the limit. These results are possible under specific properties of the approximating neural networks and due to a dedicated choice of regularization. With this, a practical contribution of this analytic paper is to provide a class of model learning frameworks different to standard settings where uniqueness can be expected in the limit of full measurements.
@@ 1 Introduction
Learning nonlinear differential equation based models from data is a highly active field of research. Its general goal is to gain information on a (partially) unknown differential-equation-based physical model from measurements of its state. Information on the model here means to either directly learn a parametrized version of the model or to learn a corresponding parametrized solution map. In both cases, neural networks are used as parametrized approximation classes in most of the existing recent works. Important examples, reviewed in [7], are physics informed neural operators [38], DeepONets [43], Fourier Neural Operators [40], Graph Neural Networks [41], Wavelet Neural Operators [56], DeepGreen [23] and model reduction [5] amongst others. In addition, we refer to the comprehensive reviews [6, 9, 13, 45] and the references therein, on the current state of the art.
@@
A natural way to overcome this, as done in [52] (and [53]) for full model learning, is to consider particular types of functions $f$: Specifying to the case $F(t,u,\varphi)=0$, a result of [52] is that a linear or algebraic function $f$ is uniquely identifiable from full state measurements if and only if the state variables (and their derivatives in case $f$ acts also on derivatives) are linearly or algebraically independent, respectively. Similarly, [52] shows that a smooth $f$ is uniquely reconstructable from full state measurements if the values of the state variables (and their derivatives) are dense in the underlying Euclidean vector space. While these results provide answers in rather general settings, the conditions on $u$ that guarantee unique recovery are difficult to verify exactly in practice ([52] provides an SVD-based algorithm that classifies unique identifiability via thresholding).
@@
where $K^{\dagger}$ is the injective full measurement operator and $y=(\hat{y}^{l})_{l}$ is the corresponding full-measurement data. With this, we allow $\mathcal{R}^{\dagger}$ to incorporate prior information on $f$ that can be used to resolve unique identifiability. In view of the above-described alternative works on uniqueness in model learning, this is related to considering $f$ to be out of a class of functions for which the measurements are sufficient for unique identifiability, only that we do not explicitly rely on such as setting, but rather provide a general framework that will always approximate a desired solution $f^{\dagger}$ as unique regularization-minimizing solution consistent with the measurement data. Of course, if the measurement data is sufficient to uniquely identify a ground-truth $f$ without the use of any additional prior information, our setting will recover this ground-truth.
In addition to the question of recovering $(\varphi^{\dagger},u^{\dagger},f^{\dagger})$ as unique solution to ($p^{\dagger}$), it is necessary to analyze in what sense parametrized solutions $(\varphi,u,f_{\theta})$ of the regularized problem
@@
While our results on using $W^{1,\infty}$ regularization are formulated for the structured-model-learning-setting here, we note that similar requirements and results can also be expected when learning learning entire PDE models e.g. with neural operators. Furthermore, regularization in general is also a standard technique to encourage loss-landscape flatness.
@@
Contributions. Following the above concept, we provide a comprehensive analysis of structured model learning in a general setting. Our main contribution is a precise mathematical setup under which we prove the above-mentioned uniqueness and approximation results. Notably, this setup differs from standard model-learning frameworks commonly used in practice, in particular with respect to the choice of regularization for the approximating functions. In view of this, a practical consequence of our work can be a suggestion of appropriate regularization functionals for model learning that ensure unique recovery in the full-measurement/ vanishing noise limit. Indeed, as an example in Appendix E shows, without appropriate regularization a unique recovery in the limit may fail.
Besides our main uniqueness result and the corresponding general framework to which it applies, we provide a well-posedness analysis and concrete examples to which our results apply. The latter includes linear and nonlinear (in the state) examples for the physical term $F$ as well as classes of neural networks for $\mathcal{F}$ to which our assumptions apply.
@@
$\partial_{t}u^{l}=\varphi^{l}\cdot\nabla u^{l}+f(u^{l})$
where $\varphi^{l}\in X_{\varphi}$ for $1\leq l\leq L$ subject to $K^{\dagger}u^{l}=\hat{y}^{l}$ with $K^{\dagger}:\mathcal{V}\to\mathcal{Y}$ an injective, linear, bounded operator and $(\hat{y}^{l})_{l}\subseteq\mathcal{Y}$ full measurement data. Assume that $f$ is approximated by neural networks $f_{\theta}$ of the form in [4, Theorem 1] parameterized by $\theta\in\Theta^{m}$ with $m\in\mathbb{N}$ a scale of approximation. Suppose that $(K^{m})_{m}$ is a sequence of bounded linear operators strongly converging to $K^{\dagger}$ and $(y^{m,l})_{m}\subseteq\mathcal{Y}$ a sequence of measurement data converging to $y^{l}$. Assume further that $U\subseteq\mathbb{R}$ is a sufficiently large interval.
@@
Furthermore, for $\lambda^{m},\mu^{m}\to\infty,\nu^{m}\to 0$ as $m\to\infty$ at certain rate depending on the neural network architectures and the measurement data $(y^{m,l})_{m}$, let $(\varphi_{m},u_{m},\theta_{m})$ be a solution to
$\min_{\varphi\in H^{1}(\Omega)^{L},u\in\mathcal{V}^{L},\theta\in\Theta^{m}}\sum_{l=1}^{L}(\|\varphi^{l}\|_{H^{1}(\Omega)}^{2}+\|u^{l}\|_{\mathcal{V}}^{2})+\|f_{\theta}\|_{L^{2}(U)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(U)}+\nu^{m}\|\theta\|$ ($p^{m}$)
$\qquad+\sum_{l=1}^{L}(\lambda^{m}\|\partial_{t}u^{l}-\varphi^{l}\cdot\nabla u^{l}-f_{\theta}(u^{l})\|^{2}_{\mathcal{W}}+\mu^{m}\|K^{m}u^{l}-y^{m,l}\|^{2}_{\mathcal{Y}}).$
@@
In contrast to the all-at-once setting pursued here, works that use a learning-informed control-to-state map to study the optimal control of certain PDEs are [15, 16, 17]. There it is assumed that the nonlinear constituents are only accessible through data-driven techniques e.g. arising from neural networks. Another related work in the field of optimal control is [12] on nonlinearity identification in the monodomain model via neural network parameterization. We also mention the recent paper [11] which deals with the identification of semilinear elliptic PDEs in a low-regularity control regime. In the context of approximating nonlinearities for elliptic state equations see [54]. We also mention the recent work [48] which establishes global convergence guarantees for adjoint-based training of infinite-width neural networks embedded in nonlinear parabolic PDEs. For structured model learning for ODEs we refer to [18, 25].
@@
For the motivation of uniqueness results for parameter identification, we refer to the works [19, 49] in the field of classical inverse problems, which derive uniqueness results based on stability estimates. Nonetheless, there is little hope to obtain results of this kind for the general system (1), even if the known physical term is linear in its physical input parameters due to the ambiguity of shift perturbations. In this respect, it seems indispensable to exploit the structural/regularity properties of the unknown term $f$ and the input parameter $\varphi$, as it is in this work and in [52], which was already discussed above. For the sake of completeness we also mention the recent preprint [30], extending the results of [52] on identifiability for symbolic recovery of differential equations to the noisy regime. Note that both works [30, 52] focus on unique identifiability per se, i.e. the classification of uniqueness, whereas our work provides an analysis-based guideline guaranteeing unique reconstructability in the limit of a practical PDE-based model learning setup.
@@ 2 Problem setting
$(t,(v_{1}^{k})_{0\leq k\leq\kappa},\dots,(v_{N}^{k})_{0\leq k\leq\kappa})$ $\mapsto f_{n}(t,(v_{1}^{k})_{0\leq k\leq\kappa},\dots,(v_{N}^{k})_{0\leq k\leq\kappa}).$
@@ 2.1 Assumptions
    for some $1\leq p,q,r,s<\infty$ with $p\geq q$, $p\geq s$. We refer to [50, Chapter 7] for the definition and properties of (Sobolev-)Bochner spaces.
@@
A possible choice of the space $V$ satisfying the compact embedding in Assumption 2 is $V=W^{\kappa+\tilde{\kappa},p_{0}}(\Omega)$ for $1<p_{0}<\infty,\tilde{\kappa}\in\mathbb{N}$ fulfilling either $\tilde{\kappa}p_{0}<d$ with $1\leq\hat{p}<\frac{dp_{0}}{d-\tilde{\kappa}p_{0}}$ or $\tilde{\kappa}p_{0}=d$ with $1\leq\hat{p}$$\infty$ due to the Rellich-Kondrachov Theorem (see e.g. [2, Theorem 6.3] and [21, §5.7]). The spaces $V_{k}$ can be chosen as $V_{k}=L^{\hat{p}}(\Omega)$ for $1\leq k\leq\kappa$.
@@
$C(0,T;W^{\tilde{\kappa},\eta}(\Omega))\hookrightarrow L^{\infty}((0,T)\times\Omega).$ (12)
@@
Note that the required convergence in (9) is rather weak in practice. In fact it holds for (potentially nonlinear) operators $(K^{m})_{m}$ converging to $K^{\dagger}$ uniformly on bounded sets in $\mathcal{V}$ (since weakly convergent sequences are bounded by [8, Proposition 3.5 (iii)]. This, in particular encompasses bounded linear operators converging in the operator norm.
@@
The assumption $\hat{f}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ in Assumption 5, iii), seems to be restrictive. However, since an admissible state $\hat{u}$ attains uniformly bounded $\mathcal{J}_{\kappa}\hat{u}$ by Assumption 5, i), the term $\hat{f}$ only acts on a compact subset of $\mathbb{R}^{D}$ in ($S$) due to its composition to $\mathcal{J}_{\kappa}\hat{u}$. Thus, whenever a $\hat{f}\in W^{1,\infty}_{loc}(\mathbb{R}^{D})^{N}$ solving ($S$) exists, there exists w.l.o.g. a solution which is globally $W^{1,\infty}(\mathbb{R}^{D})^{N}$-regular as a consequence of the following extension argument. For a sufficiently large and regular subset $U\subset\mathbb{R}^{D}$ meeting the regularity conditions in the references below and containing $(t,\mathcal{J}_{\kappa}\hat{u}(t,x))$ for a.e. $(t,x)\in(0,T)\times\Omega$ define $\hat{f}_{0}:\mathbb{R}^{D}\to\mathbb{R}^{N}$ with $\hat{f}_{0}=\hat{f}$ on $U$. The function $\hat{f}_{0}\in W^{1,\infty}(U)^{N}$ is then extendable to some $\hat{f}_{0}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ due to regularity of $U$. The result in [55, Chapter VI, Theorem 5] treats this in a more general framework that includes general Sobolev spaces and minimal smoothness conditions on the domain $U$. We refer to [2, Theorem 5.24] for an outline of the proof. A proof of the extension result, but for first-order Sobolev spaces and stricter smoothness assumptions on the underlying domain, can be found in [8, Theorem 9.7].
@@
In view of Assumption 2, i) if $V\hookrightarrow W^{\kappa+1,\hat{p}}(\Omega)$, a possible choice of the trace map $\gamma:\mathcal{V}\to\mathcal{B}$ is the (pointwise in time) Dirichlet trace operator $\gamma_{0}:V\rightarrow B$ (see [2, Chapter 5]) with $B=L^{b}(\partial\Omega)$ for $b$ as follows. Following [2, Theorem 5.36] for instance, $\gamma_{0}:W^{\kappa,\hat{p}}(\Omega)\to L^{b}(\partial\Omega)$ (and hence $\gamma$) is weak-weak continuous if $\kappa\hat{p}\leq d$ and $\hat{p}\leq b\leq\frac{(d-1)\hat{p}}{d-\kappa\hat{p}}$ (with $\hat{p}\leq b<\infty$ if $\kappa\hat{p}=d$). The choice of the (pointwise in time) Neumann trace operator (see [44, Chapter 2])) may be treated similarly with the same conditions on $b$.

### v3

@@ Abstract
This paper addresses the problem of uniqueness in learning physical laws for systems of partial differential equations (PDEs). Contrary to most existing approaches, it considers a framework of structured model learning, where existing, approximately correct physical models are augmented with components that are learned from data. The main results of the paper are a uniqueness and a convergence result that cover a large class of PDEs and a suitable class of neural networks used for approximating the unknown model components. The uniqueness result shows that, in the limit of full, noiseless measurements, a unique identification of the unknown model components as functions is possible as classical regularization-minimizing solutions of the PDE system. This result is complemented by a convergence result showing that model components learned as parameterized neural networks from incomplete, noisy measurements approximate the regularization-minimizing solutions of the PDE system in the limit. These results are possible under specific properties of the approximating neural networks and due to a dedicated choice of regularization. With this, a practical contribution of this analytic paper is to provide a class of model learning frameworks different to standard settings where uniqueness can be expected in the limit of full measurements.
@@ 1 Introduction
Learning nonlinear differential equation based models from data is a highly active field of research. Its general goal is to gain information on a (partially) unknown differential-equation-based physical model from measurements of its state. Information on the model here means to either directly learn a parametrized version of the model or to learn a corresponding parametrized solution map. In both cases, neural networks are used as parametrized approximation classes in most of the existing recent works. Important examples, reviewed in [14], are physics informed neural operators [73], DeepONets [84], Fourier Neural Operators [77], Graph Neural Networks [78], Wavelet Neural Operators [108], DeepGreen [44] and model reduction [12], among others. The comprehensive reviews [7, 13, 16, 28, 74, 89] and the references therein, provide an overview of the state of the art.
@@
A natural way to overcome this, as done in [102] (and [103]) for full model learning, is to consider particular types of functions $f$: Specifying to the case $F(t,u,\varphi)=0$, a result of [102] is that a linear or algebraic function $f$ is uniquely identifiable from full state measurements if and only if the state variables (and their derivatives in case $f$ acts also on derivatives) are linearly or algebraically independent, respectively. Similarly, [102] shows that a smooth $f$ is uniquely reconstructable from full state measurements if the values of the state variables (and their derivatives) are dense in the underlying Euclidean vector space. Consistent with this result, [104] shows that equation discovery typically requires chaotic behavior. While these results provide answers in rather general settings, the conditions on $u$ that guarantee unique recovery are difficult to verify exactly in practice ([102] provides an SVD-based algorithm that classifies unique identifiability via thresholding).
@@
where $K^{\dagger}$ is the injective full measurement operator and $y=(\hat{y}^{l})_{l}$ is the corresponding full-measurement data. With this, we allow $\mathcal{R}^{\dagger}$ to incorporate prior information on $f$ that can be used to resolve unique identifiability. In view of the above-described alternative works on uniqueness in model learning, this is related to considering $f$ to be out of a class of functions for which the measurements are sufficient for unique identifiability, only that we do not explicitly rely on such as setting, but rather provide a general framework that will always approximate a desired solution $f^{\dagger}$ as unique regularization-minimizing solution consistent with the measurement data. Of course, if the measurement data is sufficient to uniquely identify a ground-truth $f$ without the use of any additional prior information, our setting will recover this ground-truth. In addition to the question of recovering $(\varphi^{\dagger},u^{\dagger},f^{\dagger})$ as unique solution to ($p^{\dagger}$), it is necessary to analyze in what sense parametrized solutions $(\varphi,u,f_{\theta})$ of the regularized problem
@@
While our results on using $W^{1,\infty}$ regularization are formulated for the structured-model-learning-setting here, we note that similar requirements and results can also be expected when learning entire PDE models e.g. with neural operators.
@@
Contributions. Following the above concept, we provide a comprehensive analysis of structured model learning in a general setting. Our main contribution is a precise mathematical setup under which we prove the above-mentioned uniqueness and approximation results. Notably, this setup differs from standard model-learning frameworks commonly used in practice, in particular with respect to the choice of regularization for the approximating functions. In view of this, a practical consequence of our work can be a suggestion of appropriate regularization functionals for model learning that ensure unique recovery in the full-measurement/ vanishing noise limit. Indeed, as an example in Appendix E shows, without appropriate regularization a unique recovery in the limit may fail. Besides our main uniqueness result and the corresponding general framework to which it applies, we provide a well-posedness analysis and concrete examples to which our results apply. The latter includes linear and nonlinear (in the state) examples for the physical term $F$ as well as classes of neural networks for $\mathcal{F}$ to which our assumptions apply.
@@
$\partial_{t}u^{l}=\varphi^{l}\cdot\nabla u^{l}+f(u^{l})$ (4)
where $\varphi^{l}\in X_{\varphi}$ for $1\leq l\leq L$ subject to $K^{\dagger}u^{l}=\hat{y}^{l}$ with $K^{\dagger}:\mathcal{V}\to\mathcal{Y}$ an injective, linear, bounded operator and $(\hat{y}^{l})_{l}\subseteq\mathcal{Y}$ full measurement data. Suppose that there exist admissible $\hat{f}\in W^{1,\infty}(\mathbb{R})$, $\hat{u}\in\mathcal{V}^{L}$ and $\hat{\varphi}\in X_{\varphi}^{L}$ solving (4) such that $K^{\dagger}\hat{u}^{l}=\hat{y}^{l}$ for $1\leq l\leq L$. Assume that $f$ is approximated by neural networks $f_{\theta}$ of the form in [11, Theorem 1] parameterized by $\theta\in\Theta^{m}$ with $m\in\mathbb{N}$ a scale of approximation. Suppose that $(K^{m})_{m}$ is a sequence of bounded linear operators strongly converging to $K^{\dagger}$ and $(y^{m,l})_{m}\subseteq\mathcal{Y}$ a sequence of measurement data converging to $y^{l}$. Assume further that $U\subseteq\mathbb{R}$ is a sufficiently large interval.
@@
Furthermore, for $\lambda^{m},\mu^{m}\to\infty,\nu^{m}\to 0$ as $m\to\infty$ at certain rate depending on the neural network architectures and $(y^{m,l})_{m}$, let $(\varphi_{m},u_{m},\theta_{m})$ be a solution to
$\min_{\varphi\in H^{1}(\Omega)^{L},u\in\mathcal{V}^{L},\theta\in\Theta^{m}}\sum_{l=1}^{L}(\|\varphi^{l}\|_{H^{1}(\Omega)}^{2}+\|u^{l}\|_{\mathcal{V}}^{2})+\|f_{\theta}\|_{L^{2}(U)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(U)}$ ($p^{m}$)
$\qquad+\nu^{m}\|\theta\|+\sum_{l=1}^{L}(\lambda^{m}\|\partial_{t}u^{l}-\varphi^{l}\cdot\nabla u^{l}-f_{\theta}(u^{l})\|^{2}_{\mathcal{W}}+\mu^{m}\|K^{m}u^{l}-y^{m,l}\|^{2}_{\mathcal{Y}})$
@@
It is important to emphasize that, among the assumptions stated in Proposition 1, the most restrictive one from a practical perspective is the one on existence of an admissible solution $\hat{u}$ with sufficiently high regularity. This can be viewed as an implicit assumption on the existence of a sufficiently regular transport field $\hat{\varphi}$ that is compatible with $\hat{u}$. Indeed, the regularity of $\hat{u}$ as a solution to the transport equation (4) depends not only on the source term $\hat{f}$, which is essentially Lipschitz continuous, but crucially on the regularity of the transport field $\hat{\varphi}$. The relationship between the smoothness of the transport field and the well-posedness and regularity of solutions has been extensively studied. The foundational work [30] establishes well-posedness for transport equations when the transport field has Sobolev regularity, connecting the regularity of solutions to that of the flow generated by the field. A more recent survey of these results is given in [5]. Further developments, including [17, 25, 26, 27, 86], study how the smoothness of the transport field affects the stability and regularity of solutions, and demonstrate that, at critical levels of regularity, the solutions may lose uniqueness and smoothness. In fact, for Lipschitz continuous $\hat{\varphi}$ and $\hat{f}$ together with initial condition $u_{0}\in H^{2}(0,1)$ one can show, following [17, Theorem 1.2], that $u\in W^{1,2,2}(0,T;H^{1}(0,1))$. Thus, the regularity assumption on the admissible state above can be interpreted as a regularity assumption on the parameters $\hat{\varphi}$, $\hat{f}$ and $\hat{u}_{0}$.
@@
In contrast to the all-at-once setting pursued here, works that use a learning-informed control-to-state map to study the optimal control of certain PDEs are [32, 33, 34]. There it is assumed that the nonlinear constituents are only accessible through data-driven techniques e.g. arising from neural networks. Another related work in the field of optimal control is [24] on nonlinearity identification in the monodomain model via neural network parameterization. We also mention the recent paper [22] which deals with the identification of semilinear elliptic PDEs in a low-regularity control regime. In the context of approximating nonlinearities for elliptic state equations see [105]. We also mention the recent work [96] which establishes global convergence guarantees for adjoint-based training of infinite-width neural networks embedded in nonlinear parabolic PDEs. For structured model learning for ODEs we refer to [36, 46]. See also [90] on regularized inversion for hidden reaction law discovery.
Recent work incorporates conservation laws and symmetries into machine-learning models to improve physical fidelity and data efficiency. Soft-constraint methods enforce approximate conservation through regularized loss terms [57, 79, 110], whereas exact-conservation approaches leverage integral forms [51], or apply adaptive correction mechanisms [21, 43, 80]. Complementary strategies embed conservation directly into the architecture, via hard constraints in output layers [107], projection onto admissible solution spaces [93], or symmetry encoding [87], with additional architectural designs explored in [54, 81, 82, 95].
An important aspect of model learning is interpretability, which seeks representations that are accurate, parsimonious, and physically consistent. A primary goal is to recover simple laws that faithfully describe the underlying data [6, 94, 99, 101]. For a comprehensive overview of related methods, see [75].
From the perspective of inverse problems, model learning is, at its core, an identification problem, requiring that the inferred model is uniquely determined by the data to represent the true system dynamics rather than an equivalent alternative. Foundational contributions [10, 23, 31, 85] formalize structural and parameter identifiability in dynamical systems. For the motivation of uniqueness results for parameter identification, we refer to the works [20, 37, 97], which derive uniqueness from stability estimates. Uniqueness has also been established for semilinear parabolic equations [56] and for the recovery of nonlinear diffusion coefficients [67]. Further results on parameter identification for elliptic equations include [2, 4, 72]. Foundational overviews of parameter identification and related inverse problems are provided in [39] on deterministic regularization theory, [71] on coefficient inverse problems, [8] on PDE parameter estimation, and [60] on Bayesian inverse problems. Beyond coefficients, model identification targets entirely unknown PDE components, with applications to reaction-diffusion systems [35, 65, 66, 68], semilinear equations [42, 69, 70] and hyperbolic inverse sources [58, 111].
@@
Nonetheless, there is little hope to obtain results of this kind for the general system (1), even if the known physical term is linear in its physical input parameters due to the ambiguity of shift perturbations. In this respect, it seems indispensable to exploit the structural/regularity properties of the unknown term $f$ and the input parameter $\varphi$, as it is in this work and in [102], which was already discussed above. For the sake of completeness we also mention the recent preprint [52], extending the results of [102] on identifiability for symbolic recovery of differential equations to the noisy regime. Note that both works [52, 102] focus on unique identifiability per se, i.e. the classification of uniqueness, whereas our work provides an analysis-based guideline guaranteeing unique reconstructability in the limit of a practical PDE-based model learning setup.
@@ 2 Problem setting
$(t,(v_{1}^{k})_{0\leq k\leq\kappa},\dots,(v_{N}^{k})_{0\leq k\leq\kappa})$ $\mapsto f_{n}(t,(v_{1}^{k})_{0\leq k\leq\kappa},\dots,(v_{N}^{k})_{0\leq k\leq\kappa})$
@@ 2.1 Assumptions
    for some $1\leq p,q,r,s<\infty$ with $p\geq q$, $p\geq s$. We refer to [98, Chapter 7] for the definition and properties of (Sobolev-)Bochner spaces.
@@
A possible choice of the space $V$ satisfying the compact embedding in Assumption 2 is $V=W^{\kappa+\tilde{\kappa},p_{0}}(\Omega)$ for $1<p_{0}<\infty,\tilde{\kappa}\in\mathbb{N}$ fulfilling either $\tilde{\kappa}p_{0}<d$ with $1\leq\hat{p}<\frac{dp_{0}}{d-\tilde{\kappa}p_{0}}$ or $\tilde{\kappa}p_{0}=d$ with $1\leq\hat{p}$$\infty$ due to the Rellich-Kondrachov Theorem (see e.g. [3, Theorem 6.3] and [40, §5.7]). The spaces $V_{k}$ can be chosen as $V_{k}=L^{\hat{p}}(\Omega)$ for $1\leq k\leq\kappa$.
@@
$\mathcal{C}(0,T;W^{\tilde{\kappa},\eta}(\Omega))\hookrightarrow L^{\infty}((0,T)\times\Omega).$ (13)
@@
Note that the required convergence in (10) is rather weak in practice. In fact it holds for (potentially nonlinear) operators $(K^{m})_{m}$ converging to $K^{\dagger}$ uniformly on bounded sets in $\mathcal{V}$ (since weakly convergent sequences are bounded by [15, Proposition 3.5 (iii)]. This, in particular encompasses bounded linear operators converging in the operator norm.
@@
The assumption $\hat{f}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ in Assumption 5, iii), seems to be restrictive. However, since an admissible state $\hat{u}$ attains uniformly bounded $\mathcal{J}_{\kappa}\hat{u}$ by Assumption 5, i), the term $\hat{f}$ only acts on a compact subset of $\mathbb{R}^{D}$ in ($S$) due to its composition to $\mathcal{J}_{\kappa}\hat{u}$. Thus, whenever a $\hat{f}\in W^{1,\infty}_{loc}(\mathbb{R}^{D})^{N}$ solving ($S$) exists, there exists w.l.o.g. a solution which is globally $W^{1,\infty}(\mathbb{R}^{D})^{N}$-regular as a consequence of the following extension argument. For a sufficiently large and regular subset $U\subset\mathbb{R}^{D}$ meeting the regularity conditions in the references below and containing $(t,\mathcal{J}_{\kappa}\hat{u}(t,x))$ for a.e. $(t,x)\in(0,T)\times\Omega$ define $\hat{f}_{0}:\mathbb{R}^{D}\to\mathbb{R}^{N}$ with $\hat{f}_{0}=\hat{f}$ on $U$. The function $\hat{f}_{0}\in W^{1,\infty}(U)^{N}$ is then extendable to some $\hat{f}_{0}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ due to regularity of $U$. The result in [106, Chapter VI, Theorem 5] treats this in a more general framework that includes general Sobolev spaces and minimal smoothness conditions on the domain $U$. We refer to [3, Theorem 5.24] for an outline of the proof. A proof of the extension result, but for first-order Sobolev spaces and stricter smoothness assumptions on the underlying domain, can be found in [15, Theorem 9.7].
Remark 12 (Regularity of admissible state).
The existence of an admissible solution to ($S$) as required in Assumption 5, iii), with state regularity $\mathcal{V}$ may in practice be difficult to guarantee in view of the regularity typically expected from the underlying equation. For a concrete example we refer to the discussion of the transport equation following Proposition 1 in the introduction. Nevertheless, it is important to note that this regularity requirement can be interpreted as an implicit assumption on the parameter space $X_{\varphi}$, as briefly highlighted for the transport equation, since the regularity of the state is generally inherited from the regularity of the model and that of the input parameters.
@@
In view of Assumption 2, i) if $V\hookrightarrow W^{\kappa+1,\hat{p}}(\Omega)$, a possible choice of the trace map $\gamma:\mathcal{V}\to\mathcal{B}$ is the (pointwise in time) Dirichlet trace operator $\gamma_{0}:V\rightarrow B$ (see [3, Chapter 5]) with $B=L^{b}(\partial\Omega)$ for $b$ as follows. Following [3, Theorem 5.36] for instance, $\gamma_{0}:W^{\kappa,\hat{p}}(\Omega)\to L^{b}(\partial\Omega)$ (and hence $\gamma$) is weak-weak continuous if $\kappa\hat{p}\leq d$ and $\hat{p}\leq b\leq\frac{(d-1)\hat{p}}{d-\kappa\hat{p}}$ (with $\hat{p}\leq b<\infty$ if $\kappa\hat{p}=d$). The choice of the (pointwise in time) Neumann trace operator (see [88, Chapter 2])) may be treated similarly with the same conditions on $b$.